Subscribe to receive emails everytime Appbox Service Status publishes an update.
By subscribing to status updates you agree to Privacy policy and Terms of Service.
We're currently experiencing some connectivity issues with our upstream network provider.
We have contacted them and are waiting to hear back.
Update 23:42 UTC:
We're still working on the issue with them, the issue appears to be on their side.
Unfortunately we don't know what is the root cause yet.
Update 23:51 UTC:
The issue appears to be upstream at fiberring, their support have been contacted by our ISP.
We're awaiting an update.
Update 00:11 UTC:
We're still awaiting a fix from our upstream.
Update 00:22 UTC:
Our ISP are still in contact with their transit provider fiberring to fix the issue.
Update 00:43 UTC:
We're still waiting feedback from fiberring on the issue, we're still on a call with our ISP.
Update 01:01 UTC:
Our ISP is now on a call with their transit provider, we hope the issue will be resolved soon.
Update 01:26 UTC:
Fiberring confirmed that they're receiving and announcing our IP prefix, but their upstream is not.
Fiberring are now calling their carrier to resolve the issue.
Update 2:00 UTC:
We're still waiting for fiberring to resolve the issue with their upstream carrier.
Update 02:30 UTC:
Root cause has been found, fiberring's upstream require them to add us to their AS-SET.
They are instructing another on call team to update their RIPE record now.
We will hear back from them in 30 minutes.
Update 03:17 UTC:
We're still waiting for the next update, the AS-SET has not yet been updated in the RIPE DB. We continue to monitor.
Update 03:53 UTC:
The RIPE DB has been updated to add us to the AS-SET, fiberring are requesting a manual filter refresh from their upstreams.
We now have to wait for the router filters to update to the new information.
This is done automatically at certain time intervals and could potentially take hours, we're hoping if their upstreams do a manual filter refresh it will be less.
Update 04:14 UTC:
Upstreams are starting to pick up the AS-SET update, all servers are back online.